(3.3.13) The tallest living man at one time had a height of 244 cm. The shortest living man at that time had a height of 65.5 cm. Heights of men at that time had a mean of 172.29 cm and a standard deviation of 7.64 cm. Which of these two men had the height that was more​ extreme?

Since the z score for the tallest man is z=? and the z score for the shortest man is z=?​, the tallest/shortest man had the height that was more extreme
1
Expert's answer
2020-11-25T19:12:21-0500

By definition,

"\displaystyle z = \frac{x-\mu}{\sigma}"

"\displaystyle z_1 = \frac{244-172.29}{7.64}= 9.39"

"\displaystyle z_2 = \frac{65.5-172.29}{7.64}= -13.98"

The shortest man had the height that was more extreme.